# A Survey on Error-Bounded Lossy Compression for Scientific Datasets

### v1

@@ 5.8. Compression for X-ray Light Source Data
For serial crystallography, two major approaches can be combined: non-hit rejection and ROIBIN-SZ (Underwood et al., 2023). Like ptychography, the data present as 2D float-encoded integer data over a time dimension. Unlike ptychography, however, the data are substantially noisier, and there are features called Bragg spots or peaks that are key to the analysis pipeline that need to be preserved more conservatively. Non-hit rejection is a technique that uses the number of peaks detected each frame to veto or reject capturing frames that contain few, if any, peaks. While non-hit rejection eliminates on average 50% (typically between 20% and 80%) of the data from an experiment, it alone is not enough to hit the compression ratio targets for these workflows. In order to achieve higher compression ratios, a method called ROBIN-SZ was developed. ROIBIN-SZ uses the peak information used to perform non-hit rejection to losslessly preserve rectangular regions around the Bragg peaks, while using aggressive $2\times 2$ binning followed by SZ3 compression on the background. Preserving the background is critical because the peak-finding process is not infallible. There can be false-negative peaks in the dataset that need to be preserved to some extent in order for the analysis process to complete.
5.9. Compression for Data Transfer over WAN
@@
In addition to aggregating all 8 modules together to form the Ocelet framework, some other optimization strategies have been designed to improve the data transfer performance to address I/O contention, compute-node waiting, and transfer slowdown for many small files. One issue is that the compression task may exceed the capacity available on data transfer nodes or login nodes and thus require powerful compute nodes to compress the data via a batch scheduler, while such requests may not be scheduled immediately. To address this issue, Ocelet has a sentinel program to monitor and schedule the transfer/compression task dynamically. As the data transfer request is submitted, a certain amount of data will be transferred immediately without waiting for the scheduled compute resources. Whenever the compute resources are allocated by the scheduler, the remaining data that have not be transferred yet will be compressed and transferred, a process that can accelerate the overall transfer performance in turn. According to the Globus data transfer experiments across different sites over the WAN, the data transfer performance can be greatly improved by applying the adaptive parallel compression: more than 90% of the transfer time can be reduced by this method.
5.10. Compression for Boosting Communication in HPC Clusters
@@
On the collective communication front, Huang et al. (Huang et al., 2023c) developed a high-performance framework to improve performance across all MPI collectives. This CPU-based method is a significant advancement in compression-enabled MPI collective communication, demonstrating 1.8–2.7$\times$ performance improvement over traditional MPI collectives and various baselines. They also provided both theoretical analysis and experimental results to prove the limited impact of error-bounded lossy compression on the final accuracy of collective communications. However, this approach does not efficiently tackle GPU utilization, synchronization, and device-host data transfer issues, leading to suboptimal results in GPU clusters.
@@
Addressing collective communication on GPU clusters, Zhou et al. (Zhou et al., 2022) enhanced MPI_Alltoall performance on GPUs through 1D fixed-rate ZFP. Their method, however, depends on a CPU-centric staging algorithm tailored for a singular collective operation, thus limiting its applicability and performance. The fixed-rate compression further compounds these limitations, affecting both performance and compression quality. In response, Huang et al. (Huang et al., 2023b, 2024) presented a GPU-centric framework designed to optimize both collective computation and data movement, while efficiently controlling data distortion. This innovative approach harnesses the full computational capabilities of GPUs, significantly reducing the compression cost, synchronization, and device-host data transfers. The resulting performance improvements are notable, surpassing NCCL and Cray MPI by up to 4.5$\times$ and 28.7$\times$, respectively.
5.11. Compression for Distributed Machine Learning & Federated Learning Systems
@@
Recent years have witnessed the rapid evolution of deep learning models for getting high model accuracy, especially in the realm of large-scale models. Typical examples include large language foundation models (e.g., Palm and GPT-4 (Anil et al., 2023; OpenAI et al., 2023)), large-scale models in computer vision (e.g., VGGs, ResNets (Simonyan and Zisserman, 2015; He et al., 2016)), and life science (e.g., AlphaFolds (Jumper and et al., 2021)). Large-scale models have achieved significant breakthroughs and demonstrated remarkable success in learning and generative tasks in multiple domains by applying a large number of model parameters and a huge training dataset (Bommasani et al., 2022). Along with these rapid advancements and their success, however, come computational, training data collection, and privacy challenges in training these models. Distributed systems, including the public/private cloud and HPC platforms, provide strong support for training large-scale models based on large-volume training datasets to accelerate the training procedure. Therefore, distributed machine learning systems have become a hot topic in recent years. Typically, there are two distributed training schemes: data parallel (Li et al., [n. d.], 2013; Zhang et al., 2022; Zhang and Wang, 2021; Sergeev and Balso, 2018) and model parallel (Shoeybi et al., 2019).
Communication across different computing nodes during the training has been identified as the main bottleneck for distributed machine learning systems. The gradients, model parameters, and even activation data are transmitted across different computing nodes during the training. With the increase in model size, these data increase dramatically and bring high communication overhead. To improve the training performance, we need to compress them for communication reduction.
Three main kinds of communication reduction approaches exist: ❶ reducing communication rounds, ❷ communication and computing overlapping, and ❸ gradients and parameter compression. The gradient compression approaches mainly consist of gradient sparsification, gradient quantization, low rank, and error-bounded lossy compression.
@@
Gradient Sparsification The core idea here is transmitting only the gradients, which play a significant role in the model update. A gradient near zero indicates that the associated parameter has potentially been converged, and such gradients have little contribution to model updating (Zhang and Wang, 2022) and can be dropped out to reduce the communication overhead; such approaches include Top-K sparsification (Aji and Heafield, 2017; Lin et al., 2017; Sattler et al., 2019; Renggli et al., 2019). Gradient Quantization These approaches use the low-precision data to represent the original data (e.g., defined by float32 data type), which maps the discretized continuous value to different integers in a range. The one-bit SGD (Seide et al., 2014) and signSGD (Bernstein et al., 2018) open an opportunity for gradient quantization. The recent most popular gradient quantization works, TerGrad (Wen et al., 2017) and QSGD (Alistarh et al., 2017), use the stochastic unbiased estimation value of the gradient to reduce the gradient error for the gradient quantization. Low Rank Some recent research works argue that the final model learned has a “low stable rank” for the modern overparameterized DNN models (Martin and Mahoney, 2021; Li et al., 2018; Vogels et al., 2019), which can explain the impressive generalization properties of the trained DNN models. This opens an opportunity for gradient compression using the low-rank approaches, in which the large gradient matrix is decomposed as several small matrices before transferring and then reconstructed after receiving.
@@
Error-Bounded Lossy Compressors An emerging area of research is using error-bounded lossy compressors such as SZ and ZFP to reduce the size of model updates for distributed systems. Early approaches such as DeepSZ (Jin et al., 2019) focused on providing a framework for compression and storage of general DNN model architectures. Their results showed that through compressing the weights of the model, they could achieve high compression ratios in those areas, leading to reported 50$\approx$ 50$\times$ 50 × compression on the weights. Other studies, such as the FedSZ framework (Wilkins et al., 2023), demonstrate how SZ-based compression can be efficiently integrated into federated learning (FL) workflows. By compressing local model updates before transmission, FedSZ effectively reduces the bandwidth requirements and latency in FL systems, particularly in edge computing scenarios. This approach enhances communication efficiency and opens new avenues for maintaining model quality under stringent bandwidth constraints. A significant advantage of using error-bounded lossy compression is that it can retain more information than the above methods. While compressors introduce noise, they do not erase information like sparsification or quantization at certain error boundaries. Developing error-bounded lossy compression strategies to target communication reduction for distributed learning systems is an open area of research. More work is needed to address this and to understand the nuances of controlled compression algorithms and how they affect model accuracy and performance.
6. Related Work
Compression survey across domains and data types.
  • Jayasankar et al. (Jayasankar et al., 2021) contributed a comprehensive survey to summarize the data compression techniques in terms of different coding schemes (such as entropy coding and dictionary coding) amd across various data types (such as text compression, image compression, audio compression, and video compression). The survey also involved various use cases, including compression for wireless sensor networks, medical imaging, database compression, HEP data compression, and wind turbine data compression. However, the survey was written based on a high-level view of the data compression techniques, so that it has limited information about lossy compression for scientific datasets. For example, it does not mention the state-of-the-art error-bounded lossy compressors SZ (Di and Cappello, 2016; Tao et al., 2017c) and ZFP (Lindstrom, 2014) at all. Moreover, it was published in 2018, so it misses many critical lossy compression techniques such as MGARD (Ainsworth et al., 2018), SPEER (Li et al., 2023), and SZ3 (Zhao et al., 2021) developed after its publication.
  • Son et al. (Son et al., 2014) provided a survey about data compression for scientific domains that were generally produced by HPC applications. This survey involves both lossless compression techniques (such as FPC (Burtscher and Ratanaworabhan, 2009), ISOBAR (Schendel and et al., 2012), and PRIMACY (Shah et al., 2012)) and four lossy compressors (such as ISABELA (Lakshminarasimhan et al., 2011) and fpzip (Lindstrom et al., 2017)). Similar to Jayasankar et al.’s survey, however, this survey misses many key error-bounded lossy compression methods such as SZ (Di and Cappello, 2016; Tao et al., 2017c), ZFP (Lindstrom, 2014), MGARD (Ainsworth et al., 2018), SPERR (Li et al., 2023), and TTHRESH (Ballester-Ripoll et al., 2019), which were published after its publication.
Compression Survey for Specific Domains or Use Cases
  • Time series data: Chiarot et al. (Chiarot and Silvestri, 2023) provided a comprehensive survey about the principal time series compression techniques, proposing a taxonomy to classify them considering their overall approach and their characteristics. The authors also discussed the performance of the selected algorithms by comparing the experimental results that were provided in the original articles.
Compared with all these surveys, our paper has four unique features/contributions, which are summarized in next section.
@@ 7. Conclusion and Future Work
In this paper we provide a comprehensive survey to discuss error-bounded lossy compressors for scientific datasets in multiple facets. Our survey features the following key contributions:
  • We propose a lossy compression model taxonomy with 6 different compression models, from high-speed (low-quality) compression to low-speed (high-quality) compression.
  • This is the first survey paper comprehensively discussing the modular techniques commonly used in error-controlled lossy compressors.
  • This is the most comprehensive survey including 30+ state-of-the-art error-controlled lossy compressors developed from 2006 through early 2024.
  • This survey also comprehensively discusses optimized compressors customized for specific applications or use cases.
In the future, we will continue to survey modern error-bounded lossy compression techniques for diverse accelerators (such as FPGA, GPU, and CPU Vector) with various applications/use-cases.

### v2

@@ 6.9 Compression for X-ray Light Source Data
For serial crystallography, two major approaches can be combined: non-hit rejection and ROIBIN-SZ (Underwood et al., 2023b). Like ptychography, the data present as 2D float-encoded integer data over a time dimension. Unlike ptychography, however, the data are substantially noisier, and there are features called Bragg spots or peaks that are key to the analysis pipeline that need to be preserved more conservatively. Non-hit rejection is a technique that uses the number of peaks detected each frame to veto or reject capturing frames that contain few, if any, peaks. While non-hit rejection eliminates on average 50% (typically between 20% and 80%) of the data from an experiment, it alone is not enough to hit the compression ratio targets for these workflows. In order to achieve higher compression ratios, a method called ROBIN-SZ was developed. ROIBIN-SZ uses the peak information used to perform non-hit rejection to losslessly preserve rectangular regions around the Bragg peaks, while using aggressive $2\times 2$ binning followed by SZ3 compression on the background. Preserving the background is critical because the peak-finding process is not infallible. There could be false-negative peaks to preserve in the dataset in order for the analysis process to complete.
6.10 Compression for Data Transfer over WAN
@@
In addition, the Ocelet framework employs optimization strategies to enhance data transfer performance, addressing I/O contention, compute-node delays, and small-file transfer slowdowns. To avoid overloading data transfer or login nodes, Ocelet uses a sentinel program to dynamically monitor and schedule compression tasks adaptively. Initial data transfers occur immediately upon request, while remaining data is compressed and transferred once compute resources are allocated by the scheduler. This adaptive parallel compression significantly improves performance, reducing transfer times by over 90% in Globus WAN experiments (Liu et al., 2023a).
6.11 Compression for Boosting Communication in HPC Clusters
@@
On the collective communication front, Huang et al. (Huang et al., 2024c; Huang and et al., 2024a) developed CCOLL and HZCCL framework to improve performance across all MPI collectives. Such CPU-based methods are a significant advancement in compression-enabled MPI collective communication, demonstrating significant performance improvement over traditional MPI collectives and various baselines. They also provided both theoretical analysis and experimental results to prove the limited impact of error-bounded lossy compression on the final accuracy of collective communications.
@@
Addressing collective communication on GPU clusters, Zhou et al. (Zhou and et al., 2022) enhanced MPI_Alltoall performance on GPUs through 1D fixed-rate ZFP. Their method, however, depends on a CPU-centric staging algorithm tailored for a singular collective operation, thus limiting its applicability and performance. The fixed-rate compression further compounds these limitations, affecting both performance and compression quality. In response, Huang et al. (Huang et al., 2024b; Huang and et al., 2024b) presented a GPU-centric framework designed to optimize both collective computation and data movement, while efficiently controlling data distortion. This innovative approach harnesses the full computational capabilities of GPUs, significantly reducing the compression cost, synchronization, and device-host data transfers. The resulting performance improvements are notable, surpassing NCCL and Cray MPI by up to 4.5$\times$ and 28.7$\times$, respectively.
6.12 Compression for Distributed Machine Learning & Federated Learning Systems
@@
Recent years have witnessed the rapid evolution of deep learning models for getting high model accuracy, especially in the realm of large-scale models (Bommasani et al., 2022). Typical examples include large language foundation models (e.g., Palm and GPT-4 (Anil et al., 2023; OpenAI et al., 2023)), large-scale models in computer vision (e.g., VGGs, ResNets (Simonyan and Zisserman, 2015; He et al., 2016)), and life science (e.g., AlphaFolds (Jumper and et al., 2021)). Distributed systems, including the public/private cloud and HPC, provide strong support for training large-scale models based on large-volume training datasets to accelerate the training procedure (Li et al., [n. d.], 2013; Zhang et al., 2022; Zhang and Wang, 2021; sergeev2018horovods; Xu et al., 2024).
Communication between computing nodes during training has been recognized as the primary bottleneck in distributed training systems. During the training, gradients, model parameters, and activation data are transmitted across various nodes. As model sizes increase, the volume of this data escalates significantly, resulting in substantial communication overhead. To enhance training performance, it is essential to compress these data to reduce communication demands. The communication reduction approaches include: ❶ reducing communication rounds, ❷ communication and computing overlapping, and ❸ gradients and parameter compression. The gradient compression approaches mainly consist of gradient sparsification, gradient quantization, low rank, and error-bounded lossy compression.
@@
Gradient Sparsification The core idea here is transmitting only the gradients, which play a significant role in the model update (Zhang and Wang, 2022), such as Top-K sparsification (Aji and Heafield, 2017; Lin et al., 2017; Renggli et al., 2019). Gradient Quantization These approaches use the low-precision data to represent the original data (e.g., defined by float32 data type), which maps the discretized continuous value to different integers in a range, such as one-bit SGD (Seide et al., 2014), signSGD (Bernstein et al., 2018), TerGrad (Wen et al., 2017), and QSGD (Alistarh et al., 2017). Low Rank Recent works find that the learned model has a “low stable rank” for the modern overparameterized DNN models (Martin and Mahoney, 2021; Li et al., 2018; Vogels et al., 2019), which can explain the impressive generalization properties of the trained DNN models. This opens an opportunity for gradient compression using the low-rank approaches.
@@
Error-Bounded Lossy Compressors An emerging area of research is using error-bounded lossy compressors such as SZ and ZFP to reduce the size of model updates for distributed systems. Early approaches such as DeepSZ (Jin et al., 2019) focused on providing a framework for compression and storage of general DNN model architectures. Their results showed that through compressing the weights of the model, they could achieve high compression ratios in those areas, leading to reported 50$\approx$ 50$\times$ 50 × compression on the weights. Other studies, such as the FedSZ framework (Wilkins et al., 2023), demonstrate how SZ-based compression can be integrated into federated learning (FL). By compressing local model updates before transmission, FedSZ effectively reduces the bandwidth requirements and latency in FL systems, particularly in edge computing scenarios. This approach enhances communication efficiency and opens new avenues for maintaining model quality under stringent bandwidth constraints. A significant advantage of using error-bounded lossy compression is that it can retain more information than the above methods. While compressors introduce noise, they do not erase information like sparsification or quantization at certain error boundaries. Developing error-bounded lossy compression strategies to target communication reduction for distributed learning systems is an open area of research. More work is needed to address this and to understand the nuances of controlled compression algorithms and how they affect model accuracy and performance.
@@ 7. Conclusion and Future Work
In this paper we provide a comprehensive survey to discuss lossy compressors for scientific datasets in multiple facets. The key contributions include a novel compression model taxonomy, comprehensive discussion of modular compression techniques and analysis of both general-purpose scientific lossy compressors and domain-specific customized compressors.
The key takeaways from this survey include: (1) Most state-of-the-art lossy compressors have distinct strengths and weaknesses, as highlighted in our model taxonomy (Figure 1) and the overview of general-purpose compressors (Table 6), particularly in terms of throughput, compression ratios, and quality. (2) Modern lossy compressors typically rely on a combination of multiple compression techniques to achieve optimal performance, as demonstrated in Table 6, so deeply understanding each compression technique is critical to the development of effective compression algorithm. (3) Customizing lossy compressors for specific applications or use cases presents a promising avenue for development, as this can leverage the unique characteristics of specific datasets and meet user requirements can lead to highly effective compression solutions. (4) Selecting optimal compression models and approaches for diverse datasets, while tuning parameters like error bounds to meet complex user requirements for data quality, remains a challenging and active research area.
In the future, we aim to expand our survey of modern error-bounded lossy compression techniques, with a particular focus on their application across diverse devices, including emerging accelerators such as FPGAs, GPUs, and CPU vectors. Another promising avenue for exploration is the development and utilization of lossless compression techniques in scientific applications. This would complement our current survey, which primarily emphasizes lossy compression methods. In addition, most of the existing lossy compressors or compression techniques (such as prediction and wavelet transform) are depending on high autocorrelation of the data, so how to develop effective compressors for the datasets with low autocorrelations could be a future research direction.